# Short Paper

### v1

@@ Abstract
In this paper, we introduce CollaFuse, an innovative collaborative learning and inference framework designed for denoising diffusion probabilistic models. The primary objective is to address the trade-off between performance, privacy, and resource utilization—an imperative aspect for the practical implementation of information systems in real-world scenarios. Drawing inspiration from split learning [9], CollaFuse aims to balance the computationally intensive denoising process across local clients and a shared server. The framework is particularly beneficial in domains where data is scarce, private, and computational resources on local devices are limited at the client level, e.g., edge devices. This especially includes scenarios of industry [17] and healthcare [25]. CollaFuse innovatively partitions the computationally extensive denoising process into two independently trainable components. The latter remains with the client, ensuring data privacy, while the initial part is shared among clients and collaboratively learned on a server, amplifying the amount of data utilized for training and inference. Our experiment demonstrates that clients can execute numerous denoising steps on the server before client data is disclosed. The findings further indicate that a decreasing cut-ratio $c$ effectively shifts computational effort to a shared server backbone, enhancing performance generalizability. In summary, our experiment provides initial evidence supporting the advantages of collaborative learning within CollaFuse across performance, disclosed information, and local GPU energy. This highlights CollaFuse as a practical solution to the challenges of collaborative training and inference. Looking ahead, our research roadmap involves a more in-depth analysis of CollaFuse. We plan to explore the performance, encompassing both fidelity and diversity in the generated images. Additionally, our analysis includes challenging disclosed information using threat models, incorporating potential attacks like reconstruction inference. Extending our investigation, we will examine resource-related aspects, measuring floating-point operations and training durations. This thorough examination will contribute to a deeper understanding of the transformative potential of CollaFuse influencing applications in industry and healthcare and paving the way for future advancements in collaborative generative AI applications.

### v2

@@ 5 Conclusion and Outlook
In this paper, we introduce CollaFuse, an innovative collaborative learning and inference framework designed for denoising diffusion probabilistic models. The primary objective is to address the trade-off between performance, privacy, and resource utilization—an imperative aspect for the practical implementation of information systems in real-world scenarios. Drawing inspiration from split learning (gupta2018distributed), CollaFuse aims to balance the computationally intensive denoising process across local clients and a shared server. The framework is particularly beneficial in domains where data is scarce, private, and computational resources on local devices are limited. This especially includes scenarios of industry (li2022future) and healthcare (sivarajah2023responsible). CollaFuse innovatively partitions the computationally extensive denoising process into two independently trainable components. The latter remains with the client, ensuring data privacy, while the initial part is trained collaboratively on a server, amplifying the amount of training data. We demonstrate that clients can execute numerous denoising steps on the server before data is disclosed. The findings further indicate that a decreasing cut-ratio $c$ effectively shifts computational effort to a shared server backbone, enhancing performance generalizability. In summary, our experiment provides initial evidence supporting the advantages of collaborative learning within CollaFuse across performance, disclosed information, and local GPU energy. CollaFuse offers a technical solution for handling sensitive and scarce data, creating new opportunities for the IS community by exploring collaborative GenAI across fields like healthcare, edge devices, and the automotive industry. Looking ahead, our research roadmap involves a more in-depth analysis of CollaFuse. We will increase complexity by shifting to color imagery, expanding attribute diversity, and integrating text-driven generation with the Imagen model. Furthermore, we will address security by exploring attack vectors and implementing safeguards like differential privacy. This thorough examination will contribute to a deeper understanding of the transformative potential of CollaFuse influencing applications in industry and healthcare and paving the way for future advancements in collaborative generative AI applications. A promising field of research lies ahead.